Concord Inc. owns and operates a number of hardware stores in the New England region. Recently, the company has decided to locate another store in a rapidly growing area of Maryland. The company is trying to decide whether to purchase or lease the building and related facilities. Purchase: The company can purchase the site, construct the building, and purchase all store fixtures. The cost would be $1,857,100. An immediate down payment of $404,100 is required, and the remaining $1,453,000 would be paid off over 5 years at $356,900 per year (including interest payments made at end of year). The property is expected to have a useful life of 11 years, and then it will be sold for $505,800. As the owner of the property, the company will have the following out-of-pocket expenses each period. Property taxes (to be paid at the end of each year) $40,990 Insurance (to be paid at the beginning of each year) 27,040 Other (primarily maintenance which occurs at the end of each year) 17,590 $85,620 Lease: First National Bank has agreed to purchase the site, construct the building, and install the appropriate fixtures for Concord Inc. if Concord will lease the completed facility for 11 years. The annual costs for the lease would be $292,510. Concord would have no responsibility related to the facility over the 11 years. The terms of the lease are that Concord would be required to make 11 annual payments (the first payment to be made at the time the store opens and then each following year). In addition, a deposit of $91,100 is required when the store is opened. This deposit will be returned at the end of the 11th year, assuming no unusual damage to the building structure or fixtures. Click here to view factor tables Compute the present value of lease vs purchase. (Currently, the cost of funds for Concord Inc. is 10%.) (Round factor values to 5 decimal places, e.g. 1.25124 and final answer to 0 decimal places, e.g. 458,581.)

Present value   $lease $ purchase


Which of the two approaches should Concord Inc. follow?

Concord Inc. should the facilities
Answer #1

ANSWER

Present Value:

Down payment = $ 400000.

10% Present value annuity for 5 years = 3.791

10 % present value annuity for 11 years = 6.4951

10 % present value annuity due for 11 years = 7.14457

10 % Factor for 11 years = 0.3505

Cost of Buying = Down payment + Pv of Annual Installment Payment + PV of Annual property tax & other primary maintenance payment + PV of Annual Insurance Payment - pv of sale value

Down payment = 404100

Annual Installment Payment = 356900 for 5 years

Pv of Annual Installment Payment = 356900 * 3.791 = 1353008

Annual property tax & other primary maintenance payment = 40990+17590 = 58580

PV of Annual property tax & other primary maintenance payment

  =58580*6.4951=380483

Annual Insurance Payment = 27040 paid beginning of the period (use annuity due factor)

PV of Annual Insurance Payment = 27040 * 7.14457 = 193189

Sale value = 505800

pv of sale value = 505800 * 0.3505 = 177283

Cost of Buying = 404100 + 1353008 + 380483 + 193189 - 177283

Cost of Buying = 2330780 - 177283 = $2153497

.

Present Value:

Cost of Leasing = Initial Deposit + PV of annual lease payment - pv of deposit refund

Initial Deposit = 91100

Annual lease payment = 292510 ( multiply with 11year annuity due)

PV of annual lease payment = 292510 * 7.14457 = 2089858

pv of deposit refund = 100000 * 0.3505 = 35050

PV Cost of Leasing = 100000+ 2089858 - 35050

PV Cost of Leasing = $2154808

.

Lease

Buy

Present value

$2154808

$2153497

Decision : The Company should buy the building as its present value of cost purchase is lower than leasing.

================

DEAR STUDENT,

IF YOU HAVE ANY QUERY PLEASE ASK ME IN THE COMMENT BOX,I AM HERE TO HELP YOU.PLEASE GIVE ME POSITIVE RATING..

****************THANK YOU****************